OpenStudy (studen9):

The area of a rectangle knitted blanket is 15x^2-14x-8. What are the possible dimensions of the blanket. Use factoring.

OpenStudy (studen9):

ok

OpenStudy (studen9):

I think i got it

OpenStudy (studen9):

is this the answer (5x+2)(3x−4)

OpenStudy (bezerkhd):

I think you're about to get an essay. CX

imqwerty (imqwerty):

yeah thats correct factoring but its not the answer exactly we need to tell the dimensions(the length and the breadth) of the rectangle we know that area of rectangle is given by this- Area=length x breadth and the area given is this- \(15x^2-14x-8\) and we know that \(15x^2-14x-8=(5x+2)(3x-4)\) compare this^ with this-> \(area=(length)(breadth)\) compare these two equations and try to deduce the length and breadth

OpenStudy (bezerkhd):

Essay pt 2 CX

OpenStudy (bezerkhd):

Check out qwerty's new album. 2017

imqwerty (imqwerty):

yeah this factoring is also correct :) \(\large \color{red}{area}=(\color{gold}{length})(\color{blue}{breadth})\) \(\large \color{red}{15x^2-8x-4}=(\color{gold}{5x+2})(\color{blue}{3x-4})\) comparing the similar terms- \(\large \color{red}{area}=\color{red}{15x^2-8x-4}\) \(\color{gold}{length}=\color{gold}{5x+2}\) \(\color{blue}{breadth}=\color{blue}{3x-4}\)

imqwerty (imqwerty):

okay =]where are you getting confused tho?

OpenStudy (bezerkhd):

medal just cuz :D